OpenStudy (anonymous):

I will give medals and fan you!!!! Find the difference ten and two-ninths minus seven and two-thirds. Be sure to reduce to simplest form

OpenStudy (anonymous):

So you are asking what 2/9-7 and 2/3?

OpenStudy (anonymous):

yes

OpenStudy (automaticloveletter):

that's exactly what they're asking

OpenStudy (automaticloveletter):

http://www.calculatorsoup.com/calculators/math/mixednumbers.php

OpenStudy (automaticloveletter):

"= 2/9 − 23/3 = ((2 × 3) − (23 × 9)) / (9 × 3) = (6 - 207) / 27 = -201/27 = -67/9 = -7 4/9"